OpenStudy (anonymous):

You're supposed to do subsitution. -4x + y = 6 -5x - y = 21

OpenStudy (skullpatrol):

Any ideas?

hartnn (hartnn):

isolate 'y' from 1st equation by adding 4x on both sides, what you get ?

OpenStudy (skullpatrol):

Are you still there?

OpenStudy (anonymous):

Yeah.

hartnn (hartnn):

so did you add 4x to both sides ?

OpenStudy (skullpatrol):

Do you know what the "substitution principle" is?

OpenStudy (anonymous):

Aren't you supposed to substitute y first?

hartnn (hartnn):

for that you need to find y=.. from 1st equation

OpenStudy (anonymous):

@hartnn i got x + y = 6 ... Im not very bright.

OpenStudy (skullpatrol):

@aye4angela you can substitute either for "x" or "y" but you first should understand what the "substitution principle" is.

hartnn (hartnn):

-4x+y=6 Adiing 4x on both sides will give you -4x+4x+y=6+4x y=4x+6 got this ?

OpenStudy (anonymous):

Ohhhhhhhh. :3

OpenStudy (anonymous):

And then you substitute that y= into the next equation?

hartnn (hartnn):

yes.

hartnn (hartnn):

replace 'y' by '4x+6' in 2nd equation

hartnn (hartnn):

so that you get a equation in only 'x' which can be solved easily :)

hartnn (hartnn):

did you substitute that y= into the next equation ?? what you got ?

OpenStudy (skullpatrol):

You can "replace 'y' by '4x+6' in 2nd equation" because the substitution principle allows you to replace an expression by another expression that has the same value.

OpenStudy (anonymous):

On the 2nd equation i got 20x = 51. :/ i don't think im right.

hartnn (hartnn):

-5x-y =21 -5x-(4x+6)=21 how did you get 20x ? did you multiply 4 and 5 ??

OpenStudy (anonymous):

Yeah. @hartnn

hartnn (hartnn):

-5x-(4x+6)=21 now you have to open the brackets, -5x -4x -6 =21 yes! did you get this ?

OpenStudy (anonymous):

How did it turn to minus 6? That's what i always get mixed up on.

hartnn (hartnn):